OpenStudy (anonymous):

3) Which sequence is modeled by the graph below? coordinate plane showing the points (1, 2); (2, 2); (3, 2); (4, 2) A) an = 2(n − 1)1 B) an = 2(1)n − 1 C) an = 2 + (1)n − 1 D) an = (2)(n − 1) 4) Find the sum of the summation of 2 i minus 9, from i equals 3 to 10. A) 52 B) 26 C) 13 D) 32

OpenStudy (anonymous):

5) Every day, there are 3 times more likes on an internet video of a cat which is modelled by the function c(n) = (3)n − 1, where n is the number of days since the video posted. On the first day, there were 143 likes. What is the function that shows the number of likes each day? A) c(n) = (3)^(143)(n − 1) B) c(n) = (143)^n − 1 C) c(n) = (3)^143 − 1 D) c(n) = 143(3)^n − 1 6) Given the geometric sequence where a1 = −3 and the common ratio is 9, what is the domain for n? A) All integers B) All integers where n ≥ −1 C) All integers where n ≥ 1 D) All integers where n ≥ 0

OpenStudy (anonymous):

7) Grace loves to download new music. She originally had 47 songs but plans to purchase 3 new songs each week. She wants to know how many songs she will have after 104 weeks. Which equation should she use? A) a104 = 104 + 3(47 − 1) B) a104 = 104 + 3(47 + 1) C) a104 = 47 + 3(104 − 1) D) a104 = 47 + 3(104 + 1) 8) Find the sum of a finite geometric sequence from n = 1 to n = 7, using the expression −4(6)n − 1. A) 111,325 B) 526 C) 782 D) 223,948

OpenStudy (anonymous):

9) Find the 10th partial sum of the summation of 5i plus 2, from i equals 3 to infinity. A) 222 B) 240 C) 255 D) -355 10) Which sequence is modeled by the graph below? coordinate plane showing the points (1, 4); (2, 0.8); and (3, 0.16) A) an = 4(−5)n − 1 B) an = 4(one fifth)n − 1 C) an = 0.4(5)n − 1 D) an = one fifth(4)n − 1

OpenStudy (anonymous):

11) The number of visitors to a park is expected to follow the function v(x) = 5(x − 1), where x is the number of days since opening. On the first day, there will be a ceremony with 20 people in attendance. What is the function that shows total visitors, including the ceremony? A) c(x) = 20 + 5(x − 1) B) c(x) = 20 − 5(x − 1) C) c(x) = 5(x − 1) − 20 D) c(x) = 100x − 100 12) Given the geometric sequence an = 2(−3)n − 1, which of the following values for n lies in the appropriate domain for n? A) n = 1 B) n = 0 C) n = −1 D) They all lie in the domain

OpenStudy (anonymous):

did you ever find the answer to #11?